# language=Prolog, basicstyle = , moredelim = [s][](), literate = :-:-2 ,,1 ..1

### v1

@@ 2 Prior Work
Prolog [prologprogramming] is a programming language based on Horn clause logic with a long history of use in tackling problems with complex patterns. It has three types of statements: a fact is the name of a relation along with a tuple of argument values that is assumed to be part of the set defining the relation, and thus a “true” statement. A rule has as its lefthand side a similar relation name and argument tuple that is assumed to be true if all the relations/tuples in a list on the right are provably true. A query is a relation/tuple whose veracity is to be shown. The database of facts in Prolog makes it ideal to solve graph problems [Programmer-passport-prolog]. In Prolog, facts within the database can represent graph vertices, edges and properties, with the fact’s arguments denoting two vertices connected by an edge. For example, the fact edge(a,b) represents an edge between vertices a and b.
This project was inspired by the use of Prolog to encode the British Nationality Act [nat-act], where translating laws into Prolog rules allowed contradictions to be found.
@@ 3 The Graph Problem
The specific problem studied here originated from the U.S. IARPA AGILE program [agile], which provided reference implementation, data set, and timing. It is a variant of a subgraph isomorphism where the target graph is built from five types of vertices and six types of directed edges, all of which have one or more properties associated with them. Fig. 1 diagrams a simplified version of the subgraph pattern of interest [agile-workflows]. Vertices of the same type are shown by circles of the same color, labeled “1” through ”5”. Edges are also color-coded, with labels “A” through “F”; for simplicity, edges of type “A” are all edge type “F” reversed. An asterisk on a vertex or edge indicates that some property of the vertex or edge must have a specific value to satisfy the pattern  [^1] The formal description of the graph ontology has names for the vertex and edge classes, such as “person” for type 1 vertex. We use simple letters and numbers here to simplify the graph.. The dotted green lines represent a separate relationship that must hold between a property of a type A edge and that of a type 4 vertex. The red dotted line also indicates a specific relationship must hold between the two vertices of type 2. The key on the lower right gives the number of vertices for the reference dataset (about 2GB in size with about 38 million edges). For each subgraph found in the graph that matches the subgraph pattern, the type 1 vertex labelled “Root” is to be returned. A graph generator was available to create graphs of different sizes. All graphs used in this paper have exactly one matching root vertex.
@@ 4 Expression as a Logic Problem in Prolog
To express the subgraph pattern as a logic problem in Prolog, we first converted the edges and vertices of the subgraph into a database of Prolog facts, with each fact defining the properties of a vertex or edge [YinProlog]. The fact predicates are named according to the 5 vertex and 6 edge types. Furthermore, the properties of each fact are represented as Object ID in the argument values of fact. For example, a type 1 vertex is represented as the fact: vertex1(735713441679521195). This fact defines Object ID 735713441679521195 as a type 1 vertex. Similarly, edge type facts contain arguments specific to the edge. The following is the Prolog fact for a type D edge: edgeD(932362105613871012, 60, 1). The extra arguments (“1” here) correspond to specific D edge properties.
@@ 4.1 Prolog Implementation Based on Subpatterns
Prolog rules are used to express facts that depend on groups of other facts [prologprogramming]. In our first Prolog implementation, we translated each of the 6 subgraph subpatterns into separate logical rules, and a single overarching rule uses all the subpattern rules to find and return the root vertex. The clauses in the rule body are joined by conjunctions, meaning a subpattern is true if and only if all clauses are true. The goal of each subpattern rule is to identify a vertex that holds true for the all the vertex and edge properties, allowing it to be used in identifying the Root vertex. In Prolog, “variables” start with an upper case letter. For example, the following is the Prolog rule for subpattern 1:
@@
The rule states that Vertex4 is found if the following conditions are met: a D edge connects a type 4 vertex to a type 2* vertex with the Object ID 69871376, a second D edge connects the same type 4 vertex to a type 2* vertex with the Object ID 1049632, and this type 4 vertex exists in the database [^2] See [YinProlog] for the source code of the program..
@@ 6 Comparison with Conventional Implementations
  • SHAD: (Scalable High-Performance Algorithms and Data Structures) [SHAD-2018] is a C++ library that provides a common shared-memory, task-based, programming model for a variety of hardware architectures ranging from multi-core shared memory systems to multi-node distributed memory systems. The same SHAD code was run on two different systems - one with a modern hardware base but an immature runtime (denoted SHAD-1), and the second on an older and smaller system with a significantly more mature software base (denoted SHAD-2).
  • Cilk: an extension of C [10.1145/209936.209958, programming-cilk] designed to provide dynamic multi-threading in a multi-core shared memory system. Its main extensions are two keywords:
@@
  • Migrating Threads: a very Cilk-like parallel model where the underlying deeply multi-threaded hardware handles threads that must move from one physical node to another without any explicit software [kogge-refs:Dysart:2016:HSN:3018843.3018845]. This removes the need for software for both a multi-threaded runtime, and explicit software to recognize that data is “not local”. Georgia Tech’s CRNCH Center  [^5] https://research.gatech.edu/crnchs-rogues-gallery-wants-bring-weirdest-hardware-campus hosts two such systems.
@@ 7 Conclusion
Future work includes investigating the performance spike at 2 to 4 million edges, and using the built-in predicate asserta, which allows the addition of new clauses to the beginning of a database [prologprogramming]. This would reduce the number of repeated queries and decrease the number of inferences made, improving efficiency. In addition, parallel versions of Prolog may further enhance performance.

### v2

@@ 2 Prior Work
Prolog [3] is a programming language based on Horn clause logic with a long history of use in tackling problems with complex patterns. It has three types of statements: a fact is the name of a relation along with a tuple of argument values that is assumed to be part of the set defining the relation, and thus a “true” statement. A rule has as its lefthand side a similar relation name and argument tuple that is assumed to be true if all the relations/tuples in a list on the right are provably true. A query is a relation/tuple whose veracity is to be shown. The database of facts in Prolog makes it ideal to solve graph problems [9]. In Prolog, facts within the database can represent graph vertices, edges and properties, with the fact’s arguments denoting two vertices connected by an edge. For example, the fact edge(a,b) represents an edge between vertices a and b.
This project was inspired by the use of Prolog to encode the British Nationality Act [8], where translating laws into Prolog rules allowed contradictions to be found.
@@ 3 The Graph Problem
The specific problem studied here originated from the U.S. IARPA AGILE program [5], which provided reference implementation, data set, and timing. It is a variant of a subgraph isomorphism where the target graph is built from five types of vertices and six types of directed edges, all of which have one or more properties associated with them. Fig. 1 diagrams a simplified version of the subgraph pattern of interest [6]. Vertices of the same type are shown by circles of the same color, labeled “1” through ”5”. Edges are also color-coded, with labels “A” through “F”; for simplicity, edges of type “A” are all edge type “F” reversed. An asterisk on a vertex or edge indicates that some property of the vertex or edge must have a specific value to satisfy the pattern  [^1] The formal description of the graph ontology has names for the vertex and edge classes, such as “person” for type 1 vertex. We use simple letters and numbers here to simplify the graph.. The dotted green lines represent a separate relationship that must hold between a property of a type A edge and that of a type 4 vertex. The red dotted line also indicates a specific relationship must hold between the two vertices of type 2. The key on the lower right gives the number of vertices for the reference dataset (about 2GB in size with about 38 million edges). For each subgraph found in the graph that matches the subgraph pattern, the type 1 vertex labelled “Root” is to be returned. A graph generator was available to create graphs of different sizes. All graphs used in this paper have exactly one matching root vertex.
@@ 4 Expression as a Logic Problem in Prolog
To express the subgraph pattern as a logic problem in Prolog, we first converted the edges and vertices of the subgraph into a database of Prolog facts, with each fact defining the properties of a vertex or edge [10]. The fact predicates are named according to the 5 vertex and 6 edge types. Furthermore, the properties of each fact are represented as Object ID in the argument values of fact. For example, a type 1 vertex is represented as the fact: vertex1(735713441679521195). This fact defines Object ID 735713441679521195 as a type 1 vertex. Similarly, edge type facts contain arguments specific to the edge. The following is the Prolog fact for a type D edge: edgeD(932362105613871012, 60, 1). The extra arguments (“1” here) correspond to specific D edge properties.
@@ 4.1 Prolog Implementation Based on Subpatterns
Prolog rules are used to express facts that depend on groups of other facts [3]. In our first Prolog implementation, we translated each of the 6 subgraph subpatterns into separate logical rules, and a single overarching rule uses all the subpattern rules to find and return the root vertex. The clauses in the rule body are joined by conjunctions, meaning a subpattern is true if and only if all clauses are true. The goal of each subpattern rule is to identify a vertex that holds true for the all the vertex and edge properties, allowing it to be used in identifying the Root vertex. In Prolog, “variables” start with an upper case letter. For example, the following is the Prolog rule for subpattern 1:
@@
The rule states that Vertex4 is found if the following conditions are met: a D edge connects a type 4 vertex to a type 2* vertex with the Object ID 69871376, a second D edge connects the same type 4 vertex to a type 2* vertex with the Object ID 1049632, and this type 4 vertex exists in the database [^2] See [10] for the source code of the program..
@@ 6 Comparison with Conventional Implementations
  • SHAD: (Scalable High-Performance Algorithms and Data Structures) [2] is a C++ library that provides a common shared-memory, task-based, programming model for a variety of hardware architectures ranging from multi-core shared memory systems to multi-node distributed memory systems. The same SHAD code was run on two different systems - one with a modern hardware base but an immature runtime (denoted SHAD-1), and the second on an older and smaller system with a significantly more mature software base (denoted SHAD-2).
  • Cilk: an extension of C [1, 7] designed to provide dynamic multi-threading in a multi-core shared memory system. Its main extensions are two keywords:
@@
  • Migrating Threads: a very Cilk-like parallel model where the underlying deeply multi-threaded hardware handles threads that must move from one physical node to another without any explicit software [4]. This removes the need for software for both a multi-threaded runtime, and explicit software to recognize that data is “not local”. Georgia Tech’s CRNCH Center  [^5] https://research.gatech.edu/crnchs-rogues-gallery-wants-bring-weirdest-hardware-campus hosts two such systems.
@@ 7 Conclusion
Future work includes investigating the performance spike at 2 to 4 million edges, and using the built-in predicate asserta, which allows the addition of new clauses to the beginning of a database [3]. This would reduce the number of repeated queries and decrease the number of inferences made, improving efficiency. In addition, parallel versions of Prolog may further enhance performance.